# Enumerate–Conjecture–Prove: Formally Solving Answer-Construction Problems in Math Competitions

### v3

@@ 4 ConstructiveBench Dataset
In this section, we introduce the ConstructiveBench dataset, detailing its sources, curation methodology, and contents, alongside special considerations such as data contamination.
@@ 4.1 Data Sources and Scope
ConstructiveBench is curated by aggregating and rigorously filtering problems from reputable mathematical sources. Primary sources include official competition archives such as AMC 12 A/B (from 2000 onward), AIME I/II, HMMT, regional Olympiads, and the IMO Shortlist and Longlist. Additionally, we integrate data from established informal math datasets including OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024).
The dataset spans a wide difficulty spectrum, ranging from challenging high-school-level problems to sophisticated Olympiad-style questions. It covers diverse mathematical domains including algebra, number theory, combinatorics, probability, and calculus. Rather than explicit domain labeling, we employ structural filtering to retain exclusively answer-construction problems, defined as those requiring numerical or symbolic answers alongside proofs, explicitly excluding purely theorem-proving tasks without definitive solutions.
@@
Figure 3: Problem sources and problem domains in ConstructiveBench. Only the top 11 problem sources and top 7 problem domains are shown; the remaining are grouped under ‘Other’.
4.2 Data Deduplication
The dataset initially aggregated 5,555 problems. To remove duplication, we utilized both string-based matching and embedding-based duplicate filtering methods, leveraging Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic similarity threshold. This process resulted in 5,316 unique and diverse problems.
@@ 4.3 Problem Autoformalization
We used our iterative, feedback-driven autoformalization pipeline (Section 3.2) to translate each informal problem into a formally stated Lean theorem. Due to computational constraints, this process employed a combination of deepseek-v3, fine-tuned gpt-4o-mini, and fine-tuned deepseek-math-7b-instruct models, with the gpt-4o model serving as an LLM judge to verify semantic consistency.
Each problem underwent up to $T=5$ refinement iterations per model. This procedure yielded a well-formed subset of 3,431 problems, whose formalizations compile successfully in Lean, pass the LLM-based semantic check, and are formatted according to the template. Each entry in ConstructiveBench includes the informal and formal problem statements, the ground-truth answer, a formal Lean theorem, an informal solution, and metadata such as domain, source, and difficulty. Figure 4 illustrates an example dataset entry alongside the distribution of answer types present in ConstructiveBench. The diversity in answer types enables comprehensive evaluation of automated reasoning systems across various types of mathematical problems.
Furthermore, to address potential data contamination—where LLMs might inadvertently have memorized training examples—we explicitly include a carefully selected test subset of 92 problems drawn from formalized data created after June 2024, which is beyond the knowledge cutoff date of most LLMs (including those we will evaluate: gpt-4.1-nano, gpt-4.1-mini, deepseek-v3).
Figure 4: Left: Distribution of answer types in ConstructiveBench. Only the top 8 types are shown; the remaining types are grouped under ‘Other’. Right: Example dataset entry from ConstructiveBench.
@@ 5 Empirical Evaluation
The objective of our empirical evaluation was to answer the following questions.
@@
    Autoformalization Quality: Can ECP efficiently produce error-free and semantically faithful formalizations?
@@
    Answer-Construction Accuracy: Can ECP enable more accurate answer construction compared to CoT baselines?
  3. 3.
    Theorem-Proving Accuracy: Can ECP achieve higher end-to-end success rates on answer-construction problems compared to purely symbolic or LLM-only approaches?
@@ 5.1 Experimental Setup
Autoformalization.
To evaluate the quality and efficiency of our autoformalization pipeline, we measure the percentage of problems that successfully compile in Lean and are validated by our LLM judge (gpt-4o). We sequentially run up to $T=5$ iterations for three models: fine-tuned gpt-4o-mini, deepseek-math-7b-finetune, and deepseek-v3.
Answer-Construction.
For the answer construction experiment, we set the following hyperparameters: temperature = 0.0, max_tokens = 500, coder_max_attempt = 3 (until the model produces a syntactically correct Python program), and up to conjecturing_attempt = 5 iterations (until the model produces compilable Lean answers). We test light-weight commercial models including OpenAI gpt-4.1-mini and gpt-4.1-nano, as well as deepseek-v3. The models are given the formalized Lean statement and the type of answer to be constructed. We compare the ECP framework against a baseline Chain-of-Thought (CoT) method (i.e., without enumeration and pattern finding) for answer construction. The constructed answer is considered correct if and only if the equality <proposed_answer> = <ground_truth answer> can be proved by baseline symbolic methods (a sequence of proof automation tactics: simp, aesop, nlinarith, ring, norm_num).
Theorem-Proving.
We evaluate the theorem-proving tasks using state-of-the-art provers, comparing baseline symbolic methods with specialized LLM-based provers including Goedel-Prover-SFT, Kimina-Prover-7B-Preview, and DeepSeek-Prover-V2-7B. Evaluations are performed on one A100 GPU and 64 CPU cores for Lean verification, using the hyperparameters temperature=1.0, top_p=0.95, max_tokens=2048, and the Pass@32 metric. The experiment includes two conjecturing setups (CoT or ECP), three different models for conjecturing, and three different prover models for proving, resulting in a total of 18 experiments. We use 64 AMD EPYC 7453 CPU cores for 12 hours per setup, for a total of 13,824 CPU hours, and 36 GPU hours on an A100 80 GB GPU (2 hours per setup).
5.2 Results
Autoformalization.
Out of the 5,316 problems in ConstructiveBench, ECP successfully formalized 3,431 instances that were compilable and passed semantic checks, yielding an overall success rate of 64.6%. The information retriever contributed 18,516 relevant Lean definitions and theorems based on error messages, helping recover 569 additional formalizations.
Answer-Construction.
We compared the ECP framework against baseline CoT methods for answer construction on 3,431 formally verified ConstructiveBench problems and on the subset of 337 answer-construction problems from the 640 problems in PutnamBench. Answer correctness was rigorously verified through formal proofs in Lean.
The results, summarized in Table 2, clearly demonstrate ECP’s advantage over standard LLM-only methods. On the ConstructiveBench dataset, ECP substantially improved answer accuracy for gpt-4.1-mini from 14.54% (baseline) to 45.06%, gpt-4.1-nano from 17.02% to 32.59%, and deepseek-v3 from 19.06% to 43.81%. Notably, gpt-4.1-mini achieved the highest accuracy within the ECP framework.
On the uncontaminated subset of 92 problems created after June 2024, performance improvements were also pronounced: gpt-4.1-mini improved from 5.43% to 31.52%, gpt-4.1-nano from 5.43% to 20.65%, and deepseek-v3 from 4.35% to 29.35%. These significant gains demonstrate ECP’s robust generalization capabilities, particularly for novel problems beyond the training cutoff.
Model / Dataset ConstructiveBench ConstructiveBench (After Cutoff) PutnamBench
gpt-4.1-nano 17.02% / 32.59% 5.43% / 20.65% 7.72% / 11.57%
gpt-4.1-mini 14.54% / 45.06% 5.43% / 31.52% 15.13% / 23.15%
deepseek-v3 19.06% / 43.81% 4.35% / 29.35% 18.99% / 31.75%
Table 2: Answer-construction accuracy (CoT baseline / ECP) comparison on ConstructiveBench problems (including a breakdown of 92 post-knowledge-cutoff problems) and on 337 PutnamBench answer-construction problems.
Theorem-Proving.
We assess ECP’s ability to perform end-to-end formal answer construction and subsequent theorem proving using the same ConstructiveBench problems. Constructed answers generated in Section 5.2 are substituted back into the formal statements, reducing answer-construction tasks into standard theorem-proving tasks.
Prover Model
Conjecturer Model Method Symbolic Goedel-Prover-SFT Kimina-Prover-7B-Preview DeepSeek-Prover-V2-7B
gpt-4.1-nano CoT 7.41% 10.46% 11.45% 15.42%
ECP 8.11% 13.61% 16.03% 20.61%
gpt-4.1-mini CoT 6.42% 8.07% 9.24% 13.00%
ECP 9.86% 15.83% 18.51% 25.01%
deepseek-v3 CoT 6.04% 9.62% 10.49% 15.33%
ECP 9.13% 20.26% 18.27% 24.63%
Table 3: Theorem-proving accuracy (Pass@32) on 3,431 ConstructiveBench problems, comparing symbolic and LLM-based provers under CoT baseline vs. ECP-based conjecturing.
The summarized results on the ConstructiveBench problems are presented in Table 3. Among evaluated provers, DeepSeek-Prover-V2-7B consistently exhibited the best proving capability within the ECP framework, achieving the highest proving accuracy of 25.01% with gpt-4.1-mini. Goedel-Prover-SFT and Kimina-Prover-7B-Preview showed substantial improvements from baseline to ECP as well, though slightly behind DeepSeek-Prover-V2-7B. Additionally, on the subset of 337 answer-construction problems from PutnamBench, using deepseek-v3 for answer construction and DeepSeek-Prover-V2-7B for proving, ECP solves 5 problems end-to-end, achieving 1.48% accuracy, compared to 0.59% with CoT for answer construction. Although end-to-end accuracy rates are lower due to the difficulty of Putnam problems, ECP consistently enhanced proving accuracy across all model–prover combinations.
@@ 6 Limitations and Future Work
Quality of Autoformalized Theorem Statements. Our framework and dataset depend on the quality of autoformalization produced by black-box LLMs. In particular, quantitatively evaluating the quality of autoformalized statements remains a challenge in AI for mathematics. While manual evaluation offers high fidelity, it is time-consuming and difficult to scale. Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence with the original problem. Future work should explore improved evaluation metrics for formalization quality and expand datasets that better align informal and formal representations of mathematical problems.
Theorem-Proving. Our proving stage uses off-the-shelf theorem provers, which leaves significant room for optimization. In particular, designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could improve both success rates and runtime. Another limitation is the combinatorial explosion during the enumeration phase for difficult IMO problems, where the time complexity of the enumeration program grows rapidly as the in-context parameter increases. Scaling beyond enumeration demands insight and creativity akin to expert mathematicians, guiding the search toward promising conjectures rather than exhaustively exploring the entire solution space. Fine-tuning LLMs or developing search-based algorithms for the formal answer-construction task thus represents a promising avenue for next-generation mathematical AI models.
@@ 7 Conclusion
We introduced the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic approach that combines LLM-driven code enumeration, answer construction, and formal theorem proving in Lean. To support this work, we curated ConstructiveBench, a dataset of 3,431 answer-construction problems formally verified through an iterative autoformalization pipeline with error feedback. In empirical evaluations, ECP improved answer-construction accuracy with gpt-4.1-mini from 14.54% under direct CoT prompting to 45.06%. Moreover, when paired with DeepSeek-Prover-V2-7B, ECP formally proved 858 of 3,431 problems (25.01%) end-to-end. These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in solving competition-level answer-construction mathematics problems. Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to challenging Olympiad and research-level mathematical problems.
@@ Appendix B Prompt
B.1 Prompt for Formalization

### v4

@@ 4 ConstructiveBench Dataset
In this section, we introduce the ConstructiveBench dataset, detailing its sources, curation methodology, and contents, alongside special considerations such as data contamination and a human evaluation of autoformalization quality.
@@ 4.1 Data Sources and Scope
ConstructiveBench is curated by aggregating and rigorously filtering problems from reputable mathematical sources. Primary sources include official competition archives such as AMC 12 A/B (from 2000 onward), AIME I/II, HMMT, regional Olympiads, and the IMO Shortlist and Longlist. Additionally, we integrate data from established informal math datasets including OlympiadBench (He et al., 2024), Omni-Math (Gao et al., 2024a), and MathOdyssey (Fang et al., 2024). We employ structural filtering to retain exclusively answer-construction problems—those requiring numerical or symbolic answers alongside proofs—and exclude purely theorem-proving tasks without explicit answer-construction components.
4.2 Data Deduplication
The raw source dataset contains 5,555 problems. To remove duplicates, we apply both string-based matching and embedding-based filtering using Sentence-BERT (Reimers and Gurevych, 2019) and FAISS (Douze et al., 2024) with a 90% semantic-similarity threshold. This process yields 5,316 unique problems spanning a broad difficulty range—from challenging high-school problems to sophisticated Olympiad-style questions that cover diverse domains including algebra, number theory, combinatorics, probability, and calculus.
@@ 4.2 Data Deduplication
Figure 2: Problem sources and domains in ConstructiveBench. Only the top 11 sources and top 7 domains are shown; the remainder are grouped under ‘Other’.
@@ 4.3 Problem Autoformalization
To formally solve answer-construction problems in Lean, we first autoformalize each informal problem into the formal task described in Section 3.1.
Autoformalization with LLMs.
We leverage state-of-the-art autoformalization models Herald-Translator (Gao et al., 2024b) and Kimina-Autoformalizer-7B (Wang et al., 2025), together with the commercial model GPT-5 mini using a 3-shot prompt. We also use GPT-5 mini as an LLM judge. A candidate formalization is accepted if and only if both conditions hold: (i) the Lean compiler returns no errors; and (ii) the LLM judge deems the formalization semantically equivalent to the informal statement, with correct use of key definitions and no trivial restatement.
Information Retrieval in Lean.
To mitigate hallucinations and errors arising from scarce aligned informal-to-formal data (Wu et al., 2022) and rapidly evolving Lean syntax, we construct an external Lean knowledge base and an interactive loop for interaction between LLM and Lean. The knowledge base contains approximately 385K entries from the Lean 4.23.0 documentation and the mathlib repository (mathlib Community, 2020). After filtering to retain entries most relevant to high-school-level formalizations (Appendix A), the database is reduced to 57K items. Each entry is embedded with Sentence-BERT (Reimers and Gurevych, 2019) and indexed by FAISS (Douze et al., 2024); when error messages indicate missing or hallucinated definitions, we retrieve the top 5 candidates by semantic similarity and the top 5 by edit distance on the symbol name.
Error-Guided Feedback Loop.
Each of the three LLMs receives the informal problem with a 3-shot prompt and drafts a formalization. Lean compiler error messages, retrieved definitions, and LLM-judge feedback are then fed back to the model for iterative refinement, up to $T=5$ iterations. A problem is considered successfully autoformalized if any model produces an accepted formalization within the budget (a best-of-$N$ strategy across models).
Figure 3: Overview of the autoformalization pipeline: the LLM drafts a formal statement; when compilation or semantic checks fail, error messages and retrieved Lean references are fed back; the loop repeats until both checks pass.
4.4 Dataset Validation
Outputs and Coverage.
Starting from 5,316 deduplicated informal problems, we run up to $T=5$ refinement iterations per model. This produced 4,729 Lean-compilable formalizations in the correct format; among these, 3,640 also pass the LLM-judge semantic check and constituted the final ConstructiveBench dataset. Each accepted entry includes the informal statement, the formal Lean theorem, the ground-truth answer, an informal solution, and metadata such as source and difficulty. Figure 4 shows an example dataset entry and the distribution of answer types. The diversity of answer types enables comprehensive evaluation across a wide range of problem formats.
Figure 4: Left: Distribution of answer types in ConstructiveBench (top 8 shown; the remainder are grouped as ‘Other’). Right: Example dataset entry.
After-Cutoff Subset.
To mitigate contamination from pretraining corpora—where LLMs might inadvertently memorize training examples—we include a test split of 106 problems created after June 2024 (beyond the knowledge cutoff date of most LLMs we will evaluate), ensuring fair evaluation on unseen problems.
Human Evaluation.
To estimate the quality of autoformalized theorem statements in ConstructiveBench, we conducted a human evaluation on a random slice of 100 formalizations from the 3,640 accepted problems. Following the rubric in Gao et al. (2024b), each formalization was labeled as completely correct, minor error, or major error. We found 77 were fully correct, 6 had minor errors, and 17 had major errors; examples and analyses appear in Appendix C. This result aligns with human evaluations of existing autoformalized benchmarks (66.9% accuracy on Herald (Gao et al., 2024b) and 72.1% on FormalMath (Yu et al., 2025b)), highlighting the high quality of ConstructiveBench.
@@ 5 Empirical Evaluation
We evaluate whether ECP outperforms a classic chain-of-thought (CoT) baseline on two tasks:
@@
    Answer-Construction. Given a formal answer-construction problem, can the conjecturer produce a Lean expression provably equivalent to the ground truth under a fixed set of automation tactics?
@@
    End-to-End Problem Solving. Given a formal answer-construction problem, can the system both conjecture a correct answer and close the Lean goal using a prover?
@@ 5.1 Experimental Setup
Datasets.
We evaluate on ConstructiveBench (3,640 Lean-verified answer-construction problems) and the PutnamBench answer-construction subset (337 problems). We additionally report a contamination-controlled After-Cutoff split of 106 ConstructiveBench problems (created after June 2024).
Models and Hyperparameters.
For conjecturing we use GPT-5 (mini, nano), GPT-4.1 (mini, nano), and DeepSeek (V3.1, V3). For reasoning models (GPT-5 families and DeepSeek-V3.1) and lightweight non-reasoning models (GPT-4.1 families and DeepSeek-V3), we set reasoning_effort to medium and None, respectively. We use each model’s recommended decoding default parameters; for sampling we set max_tokens=4096, coder_max_attempt=3, and conjecturing_attempt=5. For proving, we use DeepSeek-Prover-V2-7B, Goedel-Prover-SFT, and Kimina-Prover-Preview-7B, with hyperparameters top_p=0.95, temperature=1.0, and max_tokens=4096. We evaluate using Pass@32 with a 120s Lean verification timeout.
Metrics.
Answer-construction accuracy marks a problem correct if the conjectured answer $a$ is provably equal to the ground truth $a_{\text{gt}}$ via the following Lean automation tactics: simp, aesop (Limperg and From, 2023), nlinarith, ring, and norm_num. End-to-end accuracy marks a problem correct if the answer is constructed by the conjecturer and the prover produces a Lean-verified proof to the theorem statement after substituting the conjectured answer. We use Pass@32 for prover, meaning that each prover model is sampled 32 outputs independently and is marked correct if at least one proof passes Lean check. Given the fixed model choice, union accuracy simply measures the fraction of successfully solved problems by any of CoT and ECP.
5.2 Answer-Construction Accuracy
In this section, we analyze the results of formal answer-construction accuracies with different conjecturer models on ConstructiveBench and PutnamBench shown in Table 2 and Table 3. Complete results for six models, including evaluations on the after-knowledge-cutoff dataset, are listed in Appendix D.
Results show that ECP consistently improves answer construction across conjecturers and datasets. On ConstructiveBench, GPT-5 mini rises from 69.7% to 73.6%, and DeepSeek-V3.1 increases from 65.4% to 75.0%, indicating that stronger conjecturers still benefit from enumeration signals. The largest absolute gains occur for lightweight non-reasoning models: DeepSeek-V3 increases from 17.3% to 40.7%. These patterns persist on the after-knowledge-cutoff split; for example, DeepSeek-V3 improves from 44.3% to 67.9%, showing robustness of ECP when data contamination is controlled.
On PutnamBench, absolute accuracies are lower due to the difficulty of the dataset, but ECP remains beneficial for all models. DeepSeek-V3.1 achieves state-of-the-art performance with 210 correct answer constructions via ECP, up from 181 problems with the CoT baseline.
In addition, analysis of union accuracy shows that CoT and ECP correctly construct answers for overlapping but distinct subsets. For example, with DeepSeek-V3.1 as conjecturer, ECP can construct correct answers for 13.8% problems that CoT baseline fails. Still, 4.2% problems fail by ECP but successful with CoT baseline because ECP’s enumeration may under-generalize or fail on several problems, misleading the conjecturer in answer construction.
Conjecturer Answer Construction End-to-End Problem Solving
CoT / ECP Union CoT / ECP Union
GPT-5 mini 69.7% / 73.6% 78.9% 32.5% / 33.1% 35.0%
DeepSeek-V3 17.3% / 40.7% 43.4% 10.1% / 18.0% 18.7%
DeepSeek-V3.1 65.4% / 75.0% 79.2% 31.1% / 32.7% 34.3%
Table 2: Answer-construction accuracy and end-to-end problem solving accuracy of CoT baseline and ECP on ConstructiveBench (3640 problems in total). The first column shows the conjecturer models for answer construction, and the second column shows the answer-construction accuracies under CoT and ECP methods, along with their union accuracy. the third column shows end-to-end accuracies using DeepSeek-Prover-V2-7B for proof generation.
Conjecturer Answer-Construction End-to-End Problem Solving
CoT / ECP Union CoT / ECP Union
GPT-5 mini 185 / 194 242 4 / 6 6
DeepSeek-V3 64 / 107 124 1 / 3 3
DeepSeek-V3.1 181 / 210 248 4 / 4 5
Table 3: Number of correct answer constructions and end-to-end solved problems on PutnamBench (337 problems in total). The first column shows the conjecturer models for answer construction, with CoT and ECP results as well as their union. The second column shows the number of problems solved end-to-end with the DeepSeek-Prover-V2-7B prover (Pass@32).
5.3 End-to-End Problem-Solving Accuracy
In this section, we discuss the results of end-to-end problem solving across various combinations of conjecturer and prover models. The rightmost columns of Table 2 and Table 3 report accuracies using different conjecturer models for answer construction, combined with DeepSeek-Prover-V2-7B for proving. The complete full per-prover tables (including Goedel-Prover-SFT and Kimina-Prover-Preview-Distill-7B) for each dataset are placed in Appendix D.
Overall, DeepSeek-Prover-V2-7B outperforms the other two prover models in these benchmarks. The strongest end-to-end problem-solving result on ConstructiveBench pairs GPT-5 mini as the conjecturer with DeepSeek-Prover-V2-7B as the prover, reaching 33.1% (up from 32.5% with CoT), with DeepSeek-V3.1 a close second at 32.7% (up from 31.1%). Even on the difficult PutnamBench answer-construction problems, ECP still solves 6 out of 337 problems end-to-end with GPT-5 mini and DeepSeek-Prover-V2-7B, compared to 4 out of 337 for CoT.
The lower overall accuracy and smaller improvements are mainly due to two factors. First, incorrectly constructed answers lead to logically wrong statements that waste prover effort. Second, the prover models have limited power and often lack the capability to formally prove difficult theorems, especially ones in PutnamBench. Overall, ECP still provides consistent end-to-end gains over CoT, with the best-performing configuration given by GPT-5 mini as conjecturer and DeepSeek-Prover-V2-7B as prover.
@@ 6 Limitations and Future Work
Quality of Autoformalized Theorem Statements. The ConstructiveBench depends on the quality of black-box LLM autoformalizations and LLM judge. In particular, quantitatively evaluating the quality of autoformalized statements remains a challenge in AI for mathematics. While manual evaluation offers high fidelity and we evaluated the autoformalization quality of a subset of 100 problems, it is still time-consuming and intractable to large scale datasets. Proxy metrics such as compile success rates or BLEU scores are more efficient but often fail to capture logical equivalence with the original problem. Future work should explore improved evaluation metrics for formalization quality and expand datasets that better align informal and formal representations of mathematical problems.
Theorem-Proving. State-of-the-art LLM is capable of winning gold medal in IMO exam (Huang and Yang, 2025), yet this achievement is out of reach for symbolic and LLM-based theorem provers in formal environment such as Lean. Our proving stage leverages off-the-shelf theorem provers, which still leaves significant room for optimization. Designing specialized domain-specific languages with inference rules tailored to high-school Olympiad mathematics could improve both success rates and runtime. Another limitation is the combinatorial explosion during the enumeration phase for difficult IMO problems, where the time complexity of the enumeration program grows rapidly as the in-context parameter increases. Scaling beyond enumeration demands insight and creativity akin to expert mathematicians, guiding the search toward promising conjectures rather than exhaustively exploring the entire solution space.
@@ 7 Conclusion
We introduced the Enumerate–Conjecture–Prove (ECP) framework, a modular neuro-symbolic approach that combines LLM-driven code enumeration, answer construction, and formal theorem proving in Lean. To support this work, we curated ConstructiveBench, a dataset of 3,640 answer‐construction problems formally verified through an iterative autoformalization pipeline with error feedback. In empirical evaluations, ECP improved the state of the art on ConstructiveBench, raising GPT-5 mini’s answer-construction accuracy from 69.7% to 73.6% and its end-to-end problem-solving accuracy from 32.5% to 33.1%. On the PutnamBench subset, ECP solves 6 out of 337 problems compared to 4 for the CoT baseline. These results demonstrate that systematically integrating enumeration and conjecture with formal verification can significantly advance the state of the art in competition-level mathematical reasoning. Future research includes enhancing autoformalization alignment and developing specialized symbolic provers tailored to Olympiad and research-level mathematical problems.
@@ Appendix B Prompt
B.1 Prompt for Autoformalization
LLM Prompt You are given a math problem that requires answer-construction. Your task is to formalize the problem in Lean 4 (v4.23.0). 1. Start with imports and namespaces such as `import Mathlib` and `open Real`. 2. Define the ground-truth answer as an abbrev: `abbrev <problem_name>_answer : $\beta$ := <expression>` or `abbrev <problem_name>_answer : $\alpha$ $\to$ $\beta$ := fun x => <expression>` (Don't put `(n : $\mathbb{N}$)` before `: $\alpha$ $\to$ $\beta$`; encode inputs inside `fun x => ...`.) 3. Theorem skeleton (leave proof as sorry): `theorem <problem_name> (x : $\alpha$) (hypotheses about x) (y : $\beta$) : P(x, y) $\leftrightarrow$ y = <problem_name>_answer x := by sorry` Other Suggestions: - Use `IsLeast` / `IsGreatest` for optimal answers. - Use `Set $\beta$` and membership if there are multiple answers.
B.2 Prompt for Enumerator
LLM Prompt You will be given a math problem that requires answer construction in Lean. Your task is to write a Python program to enumerate possible answers to assist in conjecturing and proving the true answer. For problem with unfixed context parameters or infinite answers, try several parameters and bound the number of enumerations to 100.
B.3 Prompt for Conjecturing
LLM Prompt You will be given a math problem that requires answer construction in Lean. Your task is to reason about the problem and clearly state the closed-form final answer in Lean exprssion with explanations. You do not have to derive the whole proof. The program enumeration can be used as a hint. You must not "cheat" by an answer echoing problem statement.
Appendix C Human Evaluation Examples
To verify dataset quality and estimate the false-positive error rate, we randomly sampled 100 formalizations (from the 3,640 problems which have passed both the compiler and LLM judge) for manual evaluation, labeling each as: - Correct formalization - Formalization with minor error - Formalization with major error
Below, we illustrate one example for each category. For clarity, we separate the answer and the statement, and use a placeholder for the answer variable in the Lean formalization.
1. Correct formalization: omnimath19 (Chinese Mathematical Olympiad).
Problem. Let $f:X\to X$, where $X=\{1,2,\ldots,100\}$, be a function satisfying: 1. $f(x)\neq x$ for all $x=1,2,\ldots,100$; 2. For any subset $A\subseteq X$ such that $|A|=40$, we have $A\cap f(A)\neq\emptyset$. Find the minimum $k$ such that for any such function $f$, there exists a subset $B\subseteq X$ with $|B|=k$ and $B\cup f(B)=X$. The answer is 69.
Formalization.
Listing 1: omnimath19 (correct formalization)
import Mathlib
open Set Function
abbrev omnimath19_answer : $\mathbb{N}$ := 69
theorem omnimath19 :
IsLeast {k : $\mathbb{N}$ | 0 < k $\land$
$\forall$ (f : Fin 100 $\to$ Fin 100),
($\forall$ x, f x $\neq$ x) $\to$
($\forall$ A : Finset (Fin 100), A.card = 40 $\to$ (A $\cap$ A.image f).Nonempty) $\to$
$\exists$ B : Finset (Fin 100), B.card = k $\land$ B $\cup$ B.image f = Finset.univ }
omnimath19_answer := by
sorry
Analysis. The constraint $X=\{1,\ldots,100\}$ is equivalently captured by Fin 100. The condition 1, $f(x)\neq x$, is formalized in ($\forall$ x, f x $\neq$ x). The constraint 2, $|A|=40$ and $A\cap f(A)\neq\emptyset$ is captured in using Finset.image, Finset.card and Nonempty syntax. The conditions regarding set $B$ is correctly formalized in similar manner. Finally, the target minimality condition is encoded using IsLeast. Overall, the formalization is correct without error.
2. Formalization with minor error: omnimath1410 (HMMT).
Problem. For positive integers $a,b$, let $M(a,b)=\frac{\mathrm{lcm}(a,b)}{\mathrm{gcd}(a,b)}$, and for $n\geq 2$ define
$x_{n}=M(1,M(2,M(3,\ldots,M(n-2,M(n-1,n))\ldots))).$
Compute the number of positive integers $n$ such that $2\leq n\leq 2021$ and $5x_{n}^{2}+5x_{n+1}^{2}=26x_{n}x_{n+1}$. The answer is 20.
Listing 2: omnimath1410 (formalization with minor error)
import Mathlib
open Nat
abbrev omnimath1410_answer : $\mathbb{N}$ := 20
def M (a b : $\mathbb{N}$) : $\mathbb{N}$ := Nat.lcm a b / Nat.gcd a b
def x : $\mathbb{N}$ $\to$ $\mathbb{N}$
| 0 => 0
| 1 => 1
| n + 2 => M (n + 1) (x (n + 1)) -- off-by-one here on purpose
theorem omnimath1410 :
Finset.card
(Finset.filter
(fun n => 5 * (x n)^2 + 5 * (x (n + 1))^2 = 26 * x n * x (n + 1))
(Finset.Icc 2 2021))
= omnimath1410_answer := by
sorry
Analysis. The definition of $M$ and the theorem statement are correctly formalized, but the recursive definition of $x$ is off by one. Specifically, in this formalization $x_{2}=M(1,x(1))=1$, but the intended definition gives $x_{2}=M(1,2)=2$. This can be fixed by changing M (n + 1) (x (n + 1)) to M (n + 2) (x (n + 1)).
3. Formalization with major error: IMO2023SLN7 (IMO).
Problem. Let $a,b,c,d$ be positive integers satisfying
$\frac{ab}{a+b}+\frac{cd}{c+d}=\frac{(a+b)(c+d)}{a+b+c+d}.$
Determine all possible values of $a+b+c+d$. The answer is all positive integers that are not square-free.
Listing 3: P2013AIMEII6 (formalization with major error)
import Mathlib
open Nat
noncomputable abbrev IMO2023SLN7_answer : Set $\mathbb{N}$ := { n | $\exists$ p : $\mathbb{N}$, 2 $\leq$ p $\land$ p * p | n }